OpenStudy (anonymous):

write 20% as a decimal

OpenStudy (anonymous):

20/100=1/5

OpenStudy (anonymous):

i meant 28%

OpenStudy (anonymous):

28/100=7/25

OpenStudy (anonymous):

0,28

OpenStudy (anonymous):

a decimal

OpenStudy (anonymous):

0.28 right

OpenStudy (anonymous):

0,28 is the decimal of 28/100